OpenStudy (hugy0212):

Give an example of a radical function f(x) whose domain is (3,∞) and whose range is (-∞, 0].

zepdrix (zepdrix):

Examine these four functions: \(\rm f=-\sqrt{x-3}\) \(\rm g=-\sqrt{x}-3\) \(\rm h=-\sqrt{x+3}\) \(\rm \ell=-\sqrt{x}-3\) One of these gives us the desired domain and range. I'll leave you to think about it a bit :)

OpenStudy (anonymous):

presume it means the domain is \[[3,\infty)\] right?

zepdrix (zepdrix):

Ooo I hope so :O